Reports 1-1 of 1 Clear search Modify search
Infrastructures (Air Conditioning)
soldani - 12:57 Friday 18 November 2022 (57814) Print this report
New motor anda fan for NEB AHU
The replacement of the motor-fan assembly on the NEB AHU unit was completed and the system is now working again.

The AHU was shut down from 8.30 on 17 November until 11.15 today, 18 November (local times).

The engine was restarted with the same frequency settings as the previous one, i.e. 32 Hz, drawing 8.2 amps and working at around 20% torque. The engine has been electrically connected with a "delta" configuration, as indicated on the plate, and works nominally at 1455 rpm.

(attached graph of delivery air speed measured approximately 1.5 meters after exiting the fan mouth and its pressure, before and after the works).

The dimensions of the pulleys are: motor diameter 150 mm; fan diameter mm 132

The development of the new belts is 1250 mm (net of their thickness)
d.

Images attached to this report
Comments to this report:
R. Passaquieti, D. Soldani - 17:44 Tuesday 22 November 2022 (57857) Print this report
This morning during the maintenance the fan pulley was replaced with a new one with 200 mm diameter.
Also the belt was replaced with a new one of 1407 mm length (net of the thickness).
quieti - 9:41 Thursday 24 November 2022 (57885) Print this report
Yesterday 23 Nov. at about 11:30 LT the frequency of the Inverter driving the motor of the air fan inside the NEB AHU was lowered from 32 Hz to 30 Hz.
The result of the new fan installation after the last pulley replacemmet (see prev. comment) and the rate tuning was a net reduction of the acoustic noise in the NEB Hall also in the more challanging infrasound region (see the first attached spectrogram).
Images attached to this comment
R. Passaquieti - 15:10 Thursday 24 November 2022 (57892) Print this report
I forgot to mention that yesterday the complete NEB AHU Inverter frequency tuning was from 32 Hz --> 30 Hz --> 28 Hz
Tringali, Fiori, Paoletti, Passaquieti - 18:44 Thursday 24 November 2022 (57866) Print this report

The installation of the new fan with a 200 mm pulley produced a significant acoustic broad-band noise reduction in the experimental hall, Figure 1 and Figure 2 (red curve). A seismic noise reduction is observed in narrow frequency regions, Figure 3 and Figure 4 (red curve).
The same noise reduction (blue curves) is visible in the accelerometers installed on the AHU box (*UTA*BOX), supply and return air ducts, and in the microphones of the experimental all (NN*INF_02/03/04), SAS room (*SAS_MIC and NN*INF_05) and AHU room (NN*INF_06), Figure 5 and Figure 6.
The environmental parameters of the experimental hall (pressure, temperature, humidity) are stable after the last slow-down of the inverter (28 Hz), Figure 7.
With the inverter set to 28 Hz, the frequency values of the motor-fan-belt system are:

motor ~ 13.95 Hz
fan  ~ 10.47 Hz
belt ~ 4.74 Hz

Images attached to this comment
R. Passaquieti - 17:52 Friday 25 November 2022 (57911) Print this report
This afternon at about 14:10 UTC the 200 mm diameter pulley of the air fan inside the NEB AHU was substituted with a larger one of 250 mm diameter.
The inverter frequency driving the fan rate was increased accordingly from 28 Hz to 30 Hz.
Search Help
×

Warning

×